OpenStudy (anonymous):

Anyone know how to solve this? Please help! What is the slope of the line with equation y = –3? A. –3 B. 0 C. 3 D. The slope is undefined.

OpenStudy (anonymous):

Do you know how the graph of y = -3 looks? Alternatively, ask yourself, how much does the value of y change every time the value of x changes by 1? ie, what's the value of y when x = 0? and then what's the value of y when x = 1? how much did y change by?

OpenStudy (anonymous):

To be honest with you I have problem with that, I don 't what to do. do you think I can use the formula y=mx=+c??????

OpenStudy (anonymous):

Yes you can, and m is the slope. What is the value of m if y=-3

OpenStudy (anonymous):

-3???

OpenStudy (solomonzelman):

try plugging any 2 points that lay on the line y=-3 into the SLOPE FORMULA.

OpenStudy (anonymous):

y= 0 * X + (-3)???

OpenStudy (solomonzelman):

choose any 2 points. for instance (2,-3) and (4,-3) \(\Huge\color{blue}{ \sf \frac{y_1-y_2}{x_1-x_2} }\) \(\Huge\color{blue}{ \sf \frac{(-3)-(-3)}{2-4} =? }\)

OpenStudy (anonymous):

You got it

OpenStudy (anonymous):

y = 0x - 3, so what's m?

OpenStudy (anonymous):

-3???

OpenStudy (anonymous):

No. you know that y=mx+c, so if u get that y=0x-3, that mean c is -3, so then what's m?

OpenStudy (anonymous):

I don't know how to solve it!!!

OpenStudy (anonymous):

Just compare it. m is the number which is multiplied by x, so if y = 0x -3, what;s the number that is multiplied by x

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

thank you for helping!!!

OpenStudy (anonymous):

No problem